OpenStudy (esshotwired):

What is the equation of the parabola that has a vertex of (3, 6) and a focus of (3, 8)? A) y - 6 = 1/8(x - 3)^2 B) y - 3 = 1/4(x - 6)^2 C) y - 6 = 4(x - 3)^2 D) y - 3 = 8(x - 6)^2

OpenStudy (wolfe8):

So the form y – k = a(x – h)^2 I believe. (h,k) is the vertex and you figure out to which side of the vertex is the focus. Haven't done this in awhile

OpenStudy (wolfe8):

That equation tells you which ones you can eliminate. Then use 4p(y – k) = (x – h)^2 to find p. Your p should be 2, the distance between the focus and the vertex.

OpenStudy (esshotwired):

>.< So then what do you do with p?

OpenStudy (wolfe8):

You just wanna see which equation has p=2 which corresponds to the parabola you're given

OpenStudy (esshotwired):

So how do I find which equation has p=2?

OpenStudy (wolfe8):

You rearrange the equations into the form 4p(y – k) = (x – h)^2

OpenStudy (esshotwired):

And there isn't an easier way to figure it out? >.>

OpenStudy (anonymous):

(x - h)^2 = 4p(y - k) where the vertex is (h,k) and p = distance from vertex to to focus... :|

OpenStudy (anonymous):

just put the values in it :|

OpenStudy (esshotwired):

So is the answer A?

OpenStudy (anonymous):

vertex = (3, 6) focus = (3, 8) here focus is about the vertex so p will be positive p= 2 (by distance formula) (x - 3)^2 = 4(2)(y - 6) (x - 3)^2 = 8(y - 6)

OpenStudy (anonymous):

=>1/8(x - 3)^2 = (y - 6)

OpenStudy (anonymous):

yup A

OpenStudy (esshotwired):

Okie dokie. Thanks

OpenStudy (anonymous):

:3 welcome fluffy

OpenStudy (esshotwired):

xD fluffy?

OpenStudy (anonymous):

yup your new name *-*

OpenStudy (esshotwired):

Haha ok :P